At the three consecutive corners of a square, 10 cm on each side, are point charges 4 x 10-6 C, 8 x 10-6 C, and 12 x 10-6 C, respectively. Find the EP at the fourth corner of the square.



1
Expert's answer
2022-02-06T14:32:01-0500

"q_1=4\times10^{-6}C\\q_2=8\times10^{-6}C\\q_3=12\times10^{-6}C"

"a=10cm=0.1m"

Electric field due to "q_1" Charge


"E_1=\frac{kq_1}{r^2}=\frac{9\times10^{9}\times4\times10^{-6}}{0.1^2}=36\times10^{5}N/C"

Electric field due to "q_2" charge


"E_2=\frac{kq_2}{r^2}=\frac{9\times10^{9}\times8\times10^{-6}}{(0.1\times\sqrt{2})^2}=36\times10^{5}N/C"

Electric field due to "q_3" ccharge

"E_3=\frac{kq_3}{r^2}=\frac{9\times10^{9}\times12\times10^{-6}}{(0.1^2)}=108\times10^{5}N/C"

Net electric field

"E'=\sqrt{E_1^2+E_3^2+2E_1E_3cos\theta}"


"E'=\sqrt{36^2+108^2+2\times36\times108\times cos90°}=\sqrt{(36^2+108^2)\times10^{10}}"

"E'=\sqrt{1.296\times10^{14}}=113.48\times10^5N/C"

Net electric field at point P

"E_{net}=E'+E_2"


"E_{net}=36\times10^{5}+113.48\times10^{5}=149.48\times10^5N/C"